# Bifurcations and multistability in inducible three-gene toggle switch networks

### v1

@@ V.1.1 Induction via Case (A)
To allow direct comparison to the original single-inducer model in Eqns. 2325, where $n=4$ and $\bar{a}=2$, we again set $n=4$ and now choose maximal expression to satisfy $\bar{a}+\bar{b}=2$.
@@
Fig. 13 directly compares the dynamics that emerge. Panel (A) plots the bifurcations for the baseline setting previously shown in Fig. 5, with panel (B) highlighting the fixed points at low inducer concentration in three-dimensional expression space. Panel (C) plots the bifurcations for the comparable system with self-activation in Eqns. 4648, with panel (D) again highlighting the low inducer regime. While any number of possible $(\bar{a},\bar{b})$ pairs can satisfy $\bar{a}+\bar{b}=2$, for closest comparison, we choose the limit case of $\bar{a}=2$ and $\bar{b}=0$, at which weak activator binding has a negligible impact on expression.
@@
Fig. 13(C) shows that self-activation does not affect stable expression levels or the system’s dynamics at high inducer concentrations. It does, however, influence the remaining saddle points, further accentuating expression of the dominant proteins. For example, the index-2 saddle becomes more centered in expression space at the half-maximal expression level. Also, for each index-1 saddle that includes $\bar{R}_{1}$ as a dominant protein, the dominant protein expression levels become more pronounced compared to those that are suppressed. This means that each saddle point allowing significant $\bar{R}_{1}$ expression is further from the stable points it dynamically links. If the system were to begin with half-maximal expression of each protein, for example, any trajectories leaning toward the $\bar{R}_{1}-\bar{R}_{2}$ or $\bar{R}_{1}-\bar{R}_{3}$ switches will take longer to stabilize because they must travel further to approach an index-1 saddle before veering off toward a stable point. We can thus think of these stable states as lying in deeper potential wells, making the related systems less likely to transition from one stable state to another.
@@
Figure 15: Induction cases (A) vs. (B) for competitive repressor-activator binding, with $\bar{a}=2$, $\bar{b}=2$. Each panel plots the bifurcation diagrams for $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ steady-state expressions as a function of inducer concentration. The fixed points are visualized in gene expression space below these bifurcations at low, intermediate, and high inducer concentrations. The low and high inducer concentrations chosen are the same in (A) and (B), and the intermediate inducer concentrations selected highlight the same intermediate dynamic regime. (A) The competitive repressor-activator binding case in which inducer binding controls transcription factor activity regardless of its ultimate function as a repressor or activator. (B) The competitive repressor-activator binding case in which inducer binding transforms a protein from a repressing conformation to an activating form.
@@
As one might expect, the differences between the baseline model and our chosen parametrization of the self-activation model are subtle, but the key property of self-activation is the further tunability possible through the additive expression effect from $\bar{b}$. Increasing the strength of activation while maintaining a constant maximal expression level $\bar{a}+\bar{b}$ can lead to significant dynamic transformations. Fig. 14 now fixes the system at a low inducer concentration for $\bar{a}+\bar{b}=2$, and plots the bifurcation diagram with dynamics tuned as a function of expression $\bar{b}$. As $\bar{b}$ increases from zero, all fixed points shift toward lower expression while retaining the typical dynamic profiles observed in Fig. 13. Plotting how the fixed points shown for $(\bar{a},\bar{b})=(2,0)$ would change with increasing inducer concentration, for example, would generate two dynamic phase shifts to arrive at a bistable switch between $\bar{R}_{2}$ and $\bar{R}_{3}$ expression, with results comparable to those shown in Fig. 5.
@@
With sufficiently strong activation, however, at $\bar{b}\approx 1.47$ the system no longer retains the index-2 saddle point. This is evidenced by the fixed point plot highlighted at $\bar{b}=1.5$ in Fig. 14. Moreover, $\bar{b}=1.52$ marks a bifurcation threshold from tristable to quadristable expression, where the previously index-2 fixed point essentially transforms into a stable point. Rather than comprising of a set of switches between $\bar{R}_{i}$ and $\bar{R}_{j}$ expression, the high $\bar{b}$ system now brings together a set of switches between on or off expressions for each protein. Note that, as the bifurcation point, $\bar{b}=1.52$ shares properties between the two system phases it straddles. As shown in Fig. 14, it allows the quadristable dynamics observed for higher $\bar{b}$ at low inducer concentration, but still reduces to a single bistable switch at high inducer concentration (as is characteristic of the small $\bar{b}$ regime). Systems approaching the limit where expression can only occur in the presence of bound activator, on the other hand, as in the highlighted example $(\bar{a},\bar{b})=(0,2)$ of Fig. 14, reduce down to tristability at high inducer concentration, allowing either $\bar{R}_{2}$-dominant expression, $\bar{R}_{3}$-dominant expression, or none.
V.1.2 Induction via Case (B)
@@
We now compare Case (A) above to the dynamics arising in an alternative induction setting where the effector molecule binding determines whether a protein is active as a repressor or as an activator. Fig. 12(B) outlines the thermodynamic states and weights in this setting, from which we model the system with the differential equations
@@
For reasons that become clear once plotted, we consider systems allowing a larger maximal expression of $\bar{a}+\bar{b}=4$. Specifically, we will set $\bar{a}=\bar{b}=2$ for equally strong contributions from the basal and activated expression levels.
Fig. 15 compares the system defined by induction case (A), now for $\bar{a}=\bar{b}=2$, with the corresponding model in panel (B) for the alternative induction approach highlighted in Eqns. 4951. We observe that in the low inducer concentration regime, while the dynamical structure remains the same, the saddle points allowing $\bar{R}_{1}$ expression are skewed toward favoring $\bar{R}_{1}$ expression. In this regime, a low inducer concentration means a high probability $p_{\text{act}}(c)$ of $\bar{R}_{1}$ acting as a repressor, and low probability $1-p_{\text{act}}(c)$ of acting as an activator. With $1-p_{\text{act}}(c)]\bar{R}_{1}\>\rightarrow\>0$ in Eqn. 49, $\bar{R}_{1}$ expression is essentially limited only by regulatory repression from $\bar{R}_{2}$ and $\bar{R}_{3}$, while the remaining two repressors are additionally regulated by $p_{\text{act}}(c)\bar{R}_{1}$. The resulting skewness of the central cluster to favor $\bar{R}_{1}$ places the $\bar{R}_{2}$ and $\bar{R}_{3}$-dominant stable states in deeper potential wells. It thus requires a less significant perturbation to transition out of the $\bar{R}_{1}$-dominant stable state into one of the other stable states.
Since all $\bar{R}_{1}$-expressing saddle points are spatially closer together in panel (B) compared to panel (A), this also means that case (B) of induction does not require as high an inducer concentration to transition to an intermediate dynamic regime. This also shifts the bifurcation that separates tristable from bistable dynamics to occur at a higher inducer concentration. When the inducer concentration is sufficiently high, the system transforms into one where $\bar{R}_{1}$ expression is effectively regulated by components beyond just $\bar{R}_{2}$ or $\bar{R}_{3}$. This shifts the saddle point toward a slightly higher $\bar{R}_{1}$ concentration, which serves to deepen the potential well between the $\bar{R}_{2}$ and $\bar{R}_{3}$-dominant stable states.
@@ V.2 Non-exclusive repressor-activator binding
Figure 16: Number of stable fixed points in case (A) of non-exclusive repressor-activator binding. The plots highlight stable points at low ($10^{-8}$ M) and high ($10^{-2}$ M) effector concentrations in panels (A) and (B), respectively, for varying rate parameters $\bar{a}$, $\bar{b}$, and $\bar{d}$, where $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$. The diagrams on the right hand side of panel (A) represent the types of stable points observed in each regime at low effector concentration, i.e., with eight (yellow), seven (orange), six (brown), five (green), four (teal), or three (blue) stable points. The diagrams on the right side of panel (B) represent the types of stable points observed in each regime at high effector concentration, i.e., with three (blue), two (dark blue), or one (purple) stable point(s).
@@
We now turn to systems in which repressors and activators bind non-exclusively at a gene’s regulatory region. This means that at a given promoter site, a gene’s expression can be regulated not only by the presence of one or both repressors produced by the remaining two genes, but also by the possible additional presence of bound activator. As in the competitive binding case visualized in Fig. 12, this non-exclusive binding system retains the same possible states for a given promoter site, namely (i) the state with no bound transcription factors (with an expression rate $a$), (ii) states with one or both possible repressors bound (with an expression rate 0), and (iii) states with activator bound alone (with an expression rate $a+b$). Additional states, however, account for activated expression in the presence of one or both available repressors. In these states, activated expression is suppressed by the presence of bound repressor, defined by a corresponding rate a+bda+b-d.
@@
Upon deriving a set of differential equations in dimensionless form for the expression of proteins $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ from the corresponding thermodynamic states and weights (see Appendix B for details), we have, as in the competitive binding setting, a system defined by basal expression $\bar{a}$ and additional activated expression $\bar{b}$, with the maximal expression level constrained to $\bar{a}+\bar{b}=2$ to facilitate comparison across models. Given the non-exclusive binding conditions modeled here, expression now also depends on the strength of repression from $\bar{d}$, where $\bar{d}\leq\bar{a}+\bar{b}$.
@@
The following subsections focus on the impact of effector concentration and the relative strengths of activation and repression on dynamic stability. This allows us to determine the range of dynamic landscapes possible within this parameter space under different interpretations for effector activity, and the sensitivity of the resulting dynamics and bifurcations to each of the rate parameters.
V.2.1 Induction via Case (A)
@@
We first consider Case (A) from Fig. 11, in which the binding of the effector molecule renders proteins inactive regardless of whether they are functioning as repressors or activators. Continuing our convention of choosing $\bar{R}_{1}$ as the target of effector binding, the concentration of $\bar{R}_{1}$ is thus scaled by the probability of activity $p_{\rm act}(c)$, regardless of function, where we assume that all proteins interact with effectors through the same MWC model and binding affinities. Fig. 22 and Eqns. 5961 of Appendix B define the corresponding states and weights along with the resulting dimensionless differential equations, incorporating the role of $p_{\rm act}(c)$.
@@
Figure 17: Heatmap tracking the minimum effector concentration (M) at which the stable state with no gene expression can exist for the system with non-exclusive activator-repressor binding (case (A)). Measurements span the range of rate parameters allowed by the constraints $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$, with the x-axis denoting increasing strength of repression $\bar{d}$, and the y-axis denoting decreasing strength of activation with increasing $\bar{a}$. Beige represents the parameter space in which a stable state without expression does not exist at any effector concentration.
@@
Figure 18: Heatmaps tracking the maximum effector concentrations ($M)$ at which stable states can exist with the non-exclusive activator-repressor binding model described in Fig. 11(A). The measured values represent bifurcation thresholds for the existence of these states. The heatmaps track (A) stable states expressing one of the genes that does not directly interact with effector, i.e., $(0,\bar{R}_{2},0)$ and $(0,0,\bar{R}_{3})$, (B) the stable state in which both genes not impacted by effector are expressed, i.e., $(0,\bar{R}_{2},\bar{R}_{3})$, (C) stable states expressing two genes including the effector-controlled $\bar{R}_{1}$, and (D) the stable state expressing all three genes. Measurements span the range of rate parameters allowed by the constraints $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$, with the x-axis denoting increasing strength of repression $\bar{d}$, and the y-axis denoting decreasing strength of activation with increasing $\bar{a}$. Beige represents the region of parameter space in which the stable state cannot exist at any effector concentration, and yellow indicates that the stable state is possible at all effector concentrations.
@@
Evaluating the fixed points by numerical approximation, we find that allowing non-exclusive binding expands the possible set of stable states to the maximum number of combinatorial states, i.e., no expression, one gene expressing, two genes expressing, or all three expressing. Determining the likelihood of each state being available within the allowable rate parameter space offers insight into how this more complex regulatory architecture prioritizes expression.
@@
Fig. 16(A) tracks the number of possible stable states that exist at low effector concentration ($c=10^{-8}$ M). The maximum number of stable states occurs in the yellow region where $\bar{d}\approx 0$ and $\bar{a}<0.5$ ($\bar{b}>1.5$). This represents systems in which repressor binding has negligible or no effect on expression, and gene expression relies heavily on bound activators.
@@
As $\bar{a}$ and $\bar{d}$ increase, the complexity of the dynamic landscape diminishes, with the number of possible stable points falling until $\bar{a}\gtrsim 0.5$ and $\bar{d}\gtrsim 0.5$. This threshold combination is sufficient to render the system tristable at low effector concentrations, and is similar to what we observed in the previous competitive binding models of Sections V.1.1 and V.1.2.
@@
Beyond the three possible single-gene expression states, which exist at low effector concentrations for all possible rate parameter combinations, the stable state in which both $\bar{R}_{2}$ and $\bar{R}_{3}$ dominate survives across the broadest allowable parameter space. Such a state is no longer possible only if $\bar{a}\gtrsim 0.5$ and $\bar{d}\gtrsim 0.5$. The robustness of this state makes sense given that the model in question specifically tunes $\bar{R}_{1}$ expression via effector concentration. This targeted tuning thus more indirectly (and weakly) affects the existence of a stable $(0,\bar{R}_{2},\bar{R}_{3})$ state. Fig. 16(A) also indicates that the repression parameter $\bar{d}$ drives the loss of the $(\bar{R}_{1},0,\bar{R}_{3})$ and $(\bar{R}_{1},\bar{R}_{2},0)$ states in parameter space. Essentially, once repressors play an active role in regulation (no longer $\bar{d}\rightarrow 0$), dual repressor and activator activity implies that the system can only express $\bar{R}_{1}$ significantly in a stable state when it is the only protein being expressed.
@@
Generally, as effector concentrations increase, the number of available stable fixed points decreases until the system reaches monostable, bistable, or tristable dynamics, as shown in Fig. 16(B). When $\bar{a}\lesssim 1$, such that activation contributes more strongly than basal expression, a high effector concentration suppresses all gene expression. When the system approaches the limit case of largely unregulated expression (i.e., $\bar{a}\rightarrow 2$ and $\bar{d}\rightarrow 0$) it is tristable, expressing either $\bar{R}_{2}$, $\bar{R}_{3}$, or both. Otherwise, in the relatively weak activation regime that remains, the system collapses to the bistable toggle switch between $\bar{R}_{2}$ and $\bar{R}_{3}$.
@@
Figs. 17 and 18 illustrate further how the system evolves toward these different regimes as the effector concentration increases, and how sensitive the different stable states are to a change in effector concentration. These figures plot (respectively) the minimum and maximum effector concentrations necessary for different stable fixed points to exist, where Fig. 17 specifically highlights the minimum concentration for the stable state with no gene expression. At the maximum concentration threshold, a bifurcation occurs where a sufficiently high effector concentration prevents the system from stabilizing to the state with no expression. Beige denotes the part of parameter space where a given fixed point does not exist at any effector concentration, and yellow denotes the space where the fixed point survives at high effector concentration.
@@
Notably, the heatmaps of Fig. 17 and Fig. 18(A) directly overlap, indicating a close relationship between the existence of a stable state in which no genes express, and the existence of a state that expresses one of the two genes not impacted by the effector ($\bar{R}_{2}$ or $\bar{R}_{3}$). The existence of these three stable states is thus defined almost fully by the strength of activation as reflected by the parameter $\bar{a}$. As $\bar{a}$ rises, three regimes emerge with different dynamic trends for these stable states as the effector concentration increases. When $\bar{a}\lesssim 0.5$, all three states are possible until an intermediate effector concentration is reached, beyond which only the state suppressing all expression can exist among the three. When $0\lesssim\bar{a}\lesssim 1$, we observe that a tradeoff occurs at an intermediate effector concentration, where the system’s ability to stabilize to an $\bar{R}_{2}$ or $\bar{R}_{3}$-dominant state at low concentrations is swapped for the system’s ability to suppress all expression. This indicates a regime of strong activation compared to basal expression in which an increase in effector concentration causes a fixed point to emerge and take the place of states that existed at lower concentrations. Finally, in the weak activation regime of $\bar{a}\gtrsim 1$, the system can always stabilize to an $\bar{R}_{2}$ or $\bar{R}_{3}$-dominant state at all effector concentrations and never suppresses all expression. Note in all of these regimes that the strength of repression ($\bar{d}$) has negligible or no effect on the existence of the three states.
@@
Fig. 18(B)-(D) highlights stable fixed points where more than one type of protein is expressed. The existence of such points relies on weak repression $(\bar{d}<0.5)$. As $\bar{a}$ rises and the strength of activation decreases accordingly, we observe increases in the effector concentrations at which these stable points vanish. The heatmaps indicate that as repression strength rises, a higher value of $\bar{a}$ and thus a weaker degree of activation is needed for the system to stabilize to these states, with the state in (B) responding most gradually and the state in (D) most sharply as repression strength increases. Finally, the span of effector concentrations indicates that the bifurcation thresholds for these points are more responsive to changes in activation strength than those of the single-protein expression states shown in Fig. 18(A).

### v2

@@ V.1.1 Induction in Case I
To allow direct comparison with the original single-inducer model in Eqns. 1618, we again set $n=4$ and now choose maximal expression to satisfy $\bar{a}+\bar{b}=2$.
@@
Fig. 13 compares the dynamics that emerge. Panel (A) plots the bifurcations for the baseline setting shown earlier in Fig. 5, with panel (B) highlighting the fixed points at a low inducer concentration in three-dimensional expression space. Panel (C) plots the bifurcations for the comparable system with self-activation in Eqns. 3941, with panel (D) again highlighting the low inducer regime. While any number of possible $(\bar{a},\bar{b})$ pairs can satisfy $\bar{a}+\bar{b}=2$, for closest comparison we choose the limit case of $\bar{a}=2$ and $\bar{b}=0$, where weak activator binding has a negligible impact on expression.
@@
Fig. 13(C) shows that self-activation does not affect stable expression levels or the system’s dynamics at high inducer concentrations. It does, however, influence the remaining saddle points, further accentuating expression of the dominant proteins. For example, the index-2 saddle becomes more centered in expression space at the half-maximal expression level. Also, for each index-1 saddle that favors $\bar{R}_{1}$, the dominant protein expression levels become more pronounced compared to those that are suppressed. This means that each saddle point allowing significant $\bar{R}_{1}$ expression is further from the stable points it dynamically links. If the system were to begin with half-maximal expression of each protein, for example, any trajectories leaning toward the $\bar{R}_{1}-\bar{R}_{2}$ or $\bar{R}_{1}-\bar{R}_{3}$ switches would take longer to stabilize because they must travel further to approach an index-1 saddle before veering off toward a stable point. We can think of these stable states as lying in deeper potential wells, making the related systems less likely to transition from one stable state to another.
@@
Figure 15: Induction Cases I and $I$ for competitive repressor-activator binding, with $\bar{a}=2$, $\bar{b}=2$. Each panel plots the bifurcation diagrams for $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ steady-state expressions as a function of inducer concentration. The fixed points are visualized in expression space beneath the corresponding bifurcations at low, intermediate, and high inducer concentrations. The low and high inducer concentrations chosen are the same in (A) and (B), and the intermediate concentrations selected highlight the same intermediate dynamic regime. (A) The competitive repressor-activator binding case in which inducer binding controls transcription factor activity regardless of its ultimate function as a repressor or activator (Case I). (B) The competitive repressor-activator binding case in which inducer binding transforms a protein from a repressing conformation to an activating form (Case II).
@@
As one might expect, the differences between the baseline model and our parametrization of the self-activation model are subtle, but the key advantage of self-activation lies in the additional tunability that becomes possible through the additive expression effect from $\bar{b}$. Increasing the strength of activation while maintaining a constant maximal expression level $\bar{a}+\bar{b}$ can lead to significant dynamic transformations. Fig. 14 fixes the system at a low inducer concentration for $\bar{a}+\bar{b}=2$, and plots the bifurcation diagram with dynamics tuned as a function of activation $\bar{b}$. As $\bar{b}$ increases from zero, all fixed points shift toward lower expression while retaining the typical dynamic profiles observed in Fig. 13. Plotting how the fixed points shown for $(\bar{a},\bar{b})=(2,0)$ would change with increasing inducer concentration, for example, would generate two dynamic phase shifts to arrive at a bistable switch between $\bar{R}_{2}$ and $\bar{R}_{3}$ expression, with results comparable to those shown in Fig. 5.
@@
With sufficiently strong activation at $\bar{b}\approx 1.47$, however, the system no longer retains the index-2 saddle. This is evidenced by the fixed point plot highlighted at $\bar{b}=1.5$ in Fig. 14. Moreover, $\bar{b}=1.52$ marks a threshold from tristable to quadristable expression, where the index-2 fixed point transforms into a stable point. Rather than being comprised of a set of switches between $\bar{R}_{i}$ and $\bar{R}_{j}$ expression, the high $\bar{b}$ regime now brings together a set of switches between “on” and “off” expression for each protein. Note that, as the bifurcation point, $\bar{b}=1.52$ shares properties between the two dynamic phases it straddles. As shown in Fig. 14, this point allows the quadristable dynamics observed for higher $\bar{b}$ at low inducer concentration, while still reducing to a single bistable switch at high inducer concentration (as is characteristic of the small $\bar{b}$ regime). Systems approaching the limit where expression can only occur in the presence of bound activator, on the other hand, as in the highlighted example $(\bar{a},\bar{b})=(0,2)$ of Fig. 14, will reduce down to tristability at high inducer concentration, allowing either $\bar{R}_{2}$-dominant expression, $\bar{R}_{3}$-dominant expression, or none.
V.1.2 Induction in Case II
@@
We now compare induction Case I above to the dynamics arising in an alternative setting, Case II, where the effector molecule binding determines whether a protein is active as a repressor or an activator. Fig. 12(B) outlines the thermodynamic states and weights in this setting, from which we model the system with the differential equations
@@
To aid in visual comparison with induction Case I, we consider systems allowing a larger maximal expression of $\bar{a}+\bar{b}=4$. Specifically, we will set $\bar{a}=\bar{b}=2$ for equally strong contributions from the basal and activated expression levels.
Fig. 15 compares the system defined by induction Case I, now for $\bar{a}=\bar{b}=2$, with the corresponding model in panel (B) for the alternative induction approach highlighted in Eqns. 4244. We observe that in the low inducer concentration regime, while the dynamical structure remains the same, the saddle points allowing $\bar{R}_{1}$ expression in panel (B) are now skewed toward favoring $\bar{R}_{1}$ expression. In this regime, a low inducer concentration implies a high probability $p_{\text{act}}(c)$ of $\bar{R}_{1}$ acting as a repressor, and a low probability $1-p_{\text{act}}(c)$ of acting as an activator. With $1-p_{\text{act}}(c)]\bar{R}_{1}\>\rightarrow\>0$ in Eqn. 42, $\bar{R}_{1}$ expression is limited only by regulatory repression from $\bar{R}_{2}$ and $\bar{R}_{3}$, while the remaining two repressors are additionally regulated by $p_{\text{act}}(c)\bar{R}_{1}$. The resulting skewness of the central fixed point cluster to favor $\bar{R}_{1}$ places the $\bar{R}_{2}$ and $\bar{R}_{3}$-dominant stable states in deeper potential wells. It thus requires a less significant perturbation to transition out of the $\bar{R}_{1}$-dominant stable state into one of the other stable states.
Since all $\bar{R}_{1}$-expressing saddle points are spatially closer together in the lower panel compared to the upper panel of Fig. 15, this indicates that Case II of induction does not require as high an inducer concentration to transition to an intermediate dynamic regime. Finally, at a sufficiently high inducer concentration, the saddle of the resulting bistable switch shifts the saddle point toward a slightly higher $\bar{R}_{1}$ concentration, which serves to deepen the potential well between the $\bar{R}_{2}$ and $\bar{R}_{3}$-dominant stable states compared to Case I.
@@ V.2 Non-exclusive repressor-activator binding
Figure 16: Number of stable fixed points in induction Case I of non-exclusive repressor-activator binding. The plots highlight stable points at low ($10^{-8}$ M) and high ($10^{-2}$ M) effector concentrations in panels (A) and (B), respectively, for varying rate parameters $\bar{a}$, $\bar{b}$, and $\bar{d}$, where $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$. The diagrams on the right hand side of panel (A) represent the types of stable points observed in each regime at a low effector concentration, i.e., with eight (yellow), seven (orange), six (brown), five (green), four (teal), or three (blue) stable points. The diagrams on the right side of panel (B) represent the types of stable points observed in each regime at a high effector concentration, i.e., with three (blue), two (dark blue), or one (purple) stable point(s).
@@
We now turn to systems in which repressors and activators bind non-exclusively to a gene’s regulatory region. This means that at a given promoter site, a gene’s expression can be regulated not only by the presence of one or both repressors produced by the remaining two genes, but also by the possible additional presence of bound activator. As in the competitive binding case shown in Fig. 12, this system with non-exclusive binding retains the same possible states for a given promoter site, namely (i) the state with no bound transcription factors (with an expression rate $a$), (ii) states with one or both possible repressors bound (with an expression rate 0), and (iii) states with activator bound alone (with an expression rate $a+b$). Additional states account for activated expression in the presence of one or both available repressors, and in these states, activated expression is suppressed by the presence of bound repressor defined by a corresponding rate a+bda+b-d.
@@
Upon deriving a set of differential equations in dimensionless form for the expression of proteins $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ from the corresponding thermodynamic states and weights (see Appendix C for details), we have, as in the competitive binding setting, a system defined by basal expression $\bar{a}$ and additional activated expression $\bar{b}$, with the maximal expression level constrained once again to $\bar{a}+\bar{b}=2$ to facilitate comparison across models. Given the non-exclusive binding conditions modeled, expression now also depends on the strength of repression from $\bar{d}$, where $\bar{d}\leq\bar{a}+\bar{b}$.
@@
The following subsections focus on the impact of effector concentration and the relative strengths of activation and repression on dynamic stability. This allows us to determine the range of dynamic landscapes possible within the parameter space under different interpretations for effector activity, and the sensitivity of the resulting dynamics and bifurcations to each of the rate parameters.
V.2.1 Induction in Case I
@@
We first consider Case I from Fig. 11(A), in which the binding of the effector molecule renders proteins inactive regardless of whether they are functioning as repressors or activators. Continuing our convention of choosing $\bar{R}_{1}$ as the target of effector binding, the concentration of $\bar{R}_{1}$ is scaled by the probability of activity $p_{\rm act}(c)$, regardless of function, where we assume that all proteins interact with effectors through the same MWC model and binding affinities. Fig. 22 and Eqns. 5961 of Appendix C define the corresponding states and weights along with the resulting dimensionless differential equations incorporating the role of $p_{\rm act}(c)$.
@@
Figure 17: Heatmap tracking the minimum effector concentration (M) at which the stable state with no expression can exist with non-exclusive activator-repressor binding (induction Case I). Measurements span the range of rate parameters allowed by the constraints $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$, with the x-axis denoting increasing strength of repression $\bar{d}$, and the y-axis denoting decreasing strength of activation with increasing $\bar{a}$. The beige region represents the parameter space in which a stable state without expression does not exist at any effector concentration.
@@
Figure 18: Heatmaps tracking the maximum effector concentrations ($M)$ (bifurcation thresholds) at which stable states can exist for the non-exclusive activator-repressor binding model described in Fig. 11(A) (Case I). The heatmaps track (A) stable states expressing one of the genes that does not directly interact with an effector, i.e., $(0,\bar{R}_{2},0)$ and $(0,0,\bar{R}_{3})$, (B) the stable state in which the two genes not impacted by effector are expressed, i.e., $(0,\bar{R}_{2},\bar{R}_{3})$, (C) stable states expressing two genes including the effector-controlled $\bar{R}_{1}$, and (D) the stable state expressing all three genes. Measurements span the rate parameters allowed by the constraints $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$, with the x-axis denoting increasing repression strength $\bar{d}$, and the y-axis denoting decreasing activation with increasing $\bar{a}$. Beige represents the region of parameter space in which the stable state cannot exist at any effector concentration, and the yellow-shaded region indicates where the stable state is possible at all effector concentrations.
@@
Evaluating the fixed points, we find that allowing non-exclusive binding expands the possible set of stable states to the maximum combinatorial number, i.e., no expression, one gene expressing, two genes expressing, or all three expressing. Determining the likelihood of each state being accessible within the allowable rate parameter space offers insights into how this more complex regulatory architecture prioritizes expression.
@@
Fig. 16(A) tracks the number of possible stable states that exist at low effector concentration ($c=10^{-8}$ M). The maximum number of stable states occurs in the yellow-shaded region where $\bar{d}\approx 0$ and $\bar{a}<0.5$ ($\bar{b}>1.5$). This represents systems in which repressor binding has negligible effects on expression, and where gene expression thus relies heavily on bound activators.
@@
As $\bar{a}$ and $\bar{d}$ rise, the complexity of the dynamic landscape decreases, with the number of possible stable points falling until $\bar{a}\gtrsim 0.5$ and $\bar{d}\gtrsim 0.5$. This threshold combination is sufficient to render the system tristable at low effector concentrations, and is similar to what we observed in the previous competitive binding models of Sections V.1.1 and V.1.2.
@@
Other than the three possible single-gene expression states, which exist at low effector concentrations for all possible rate parameter combinations, the stable state that survives across the broadest allowable parameter space is the one in which both $\bar{R}_{2}$ and $\bar{R}_{3}$ dominate. Such a state becomes no longer possible only if $\bar{a}\gtrsim 0.5$ and $\bar{d}\gtrsim 0.5$. The robustness of this state makes sense given that the model in question specifically tunes $\bar{R}_{1}$ expression via effector concentration. This targeted tuning more indirectly (and weakly) affects the existence of a stable $(0,\bar{R}_{2},\bar{R}_{3})$ state. Fig. 16(A) also indicates that the repression parameter $\bar{d}$ drives the loss of the $(\bar{R}_{1},0,\bar{R}_{3})$ and $(\bar{R}_{1},\bar{R}_{2},0)$ states. Essentially, once repressors play an active role in regulation (no longer $\bar{d}\rightarrow 0$), dual repressor and activator activity implies that the system can only express $\bar{R}_{1}$ significantly in a stable state when it is the only protein being expressed.
@@
Generally, as effector concentrations increase, the number of available stable fixed points decreases until the system reaches monostable, bistable, or tristable dynamics, as shown in Fig. 16(B). When $\bar{a}\lesssim 1$, such that activation contributes more strongly than basal expression, a high effector concentration suppresses all gene expression. When the system approaches the limit case of largely unregulated expression (i.e., $\bar{a}\rightarrow 2$ and $\bar{d}\rightarrow 0$) it becomes tristable, expressing either $\bar{R}_{2}$, $\bar{R}_{3}$, or both. Otherwise, in the relatively weak activation regime that remains, the system collapses to the bistable toggle switch between $\bar{R}_{2}$ and $\bar{R}_{3}$.
@@
Figs. 17 and 18 illustrate further how the system evolves toward these different regimes as the effector concentration rises, and how sensitive the different stable states are to these changes. The figures plot (respectively) the minimum and maximum effector concentrations that are required for different stable fixed points to exist, where Fig. 17 specifically highlights the minimum concentration for the stable state with no expression. At the maximum concentration threshold, a bifurcation occurs when a sufficiently high effector concentration prevents the system from stabilizing to the state with no expression. The beige-shaded region denotes the part of parameter space where a given fixed point does not exist at any effector concentration, and the yellow-shaded region denotes the space where the fixed point survives at high effector concentration.
@@
Notably, the heatmaps of Fig. 17 and Fig. 18(A) directly overlap, indicating a close relationship between the existence of a stable state with no expression and the existence of a state that expresses one of the two proteins not impacted by the effector ($\bar{R}_{2}$ or $\bar{R}_{3}$). The existence of these three stable states is thus defined almost fully by the strength of activation as reflected by $\bar{a}$. As $\bar{a}$ rises, three regimes with distinct dynamic trends emerge upon tuning the effector concentration. When $\bar{a}\lesssim 0.5$, all three states are possible until an intermediate effector concentration is reached, beyond which only the state suppressing all expression can exist among the three. When $0\lesssim\bar{a}\lesssim 1$, we observe a tradeoff at an intermediate effector concentration, where the system’s ability to stabilize to an $\bar{R}_{2}$ or $\bar{R}_{3}$-dominant state at low concentrations is swapped for the system’s ability to suppress all expression. This indicates a regime of strong activation compared to basal expression, where an increase in effector concentration causes a fixed point to emerge and replace states that existed at lower concentrations. Finally, in the weak activation regime of $\bar{a}\gtrsim 1$, the system can always stabilize to an $\bar{R}_{2}$ or $\bar{R}_{3}$-dominant state at all effector concentrations, and never suppresses expression entirely. Note in all of these regimes that the strength of repression ($\bar{d}$) has negligible or no effect on the existence of the three states.
@@
Fig. 18(B)-(D) highlights stable fixed points where more than one type of protein is expressed. The existence of such points relies on weak repression $(\bar{d}<0.5)$. As $\bar{a}$ rises and the strength of activation decreases accordingly, we observe levels of rising effector concentrations where the stable points vanish. The heatmaps indicate that as repression strength rises, a higher value of $\bar{a}$ and thus a weaker degree of activation is needed for the system to stabilize to these states, with the state in panel (B) responding most gradually and the state in panel (D) most sharply as repression strength increases. Finally, the span of effector concentrations indicates that the bifurcation thresholds for these stable points are more responsive to changes in activation strength than those of the single-protein expression states shown in Fig. 18(A).